# Object-Centric Image to Video Generation with Language Guidance

### v1

@@ Abstract
Accurate and flexible world models are crucial for autonomous systems to understand their environment and predict future events. Object-centric models, with structured latent spaces, have shown promise in modeling object dynamics and interactions, but often face challenges in scaling to complex datasets and incorporating external guidance, limiting their applicability in robotics. To address these limitations, we propose TextOCVP, an object-centric model for image-to-video generation guided by textual descriptions. TextOCVP parses an observed scene into object representations, called slots, and utilizes a text-conditioned transformer predictor to forecast future object states and video frames. Our approach jointly models object dynamics and interactions while incorporating textual guidance, thus leading to accurate and controllable predictions. Our method’s structured latent space offers enhanced control over the prediction process, outperforming several image-to-video generative baselines. Additionally, we demonstrate that structured object-centric representations provide superior controllability and interpretability, facilitating the modeling of object dynamics and enabling more precise and understandable predictions. Videos and code are available at https://play-slot.github.io/TextOCVP/.
Object-Centric Learning, Text-Conditioned Video Prediction, Object-Centric Video Prediction, Representation Learning, Transformers
Refer to captiona) Object-Centric DecompositionLearned ObjectsRefer to captionb) Text-Conditioned Image-to-Video Generation
@@
$\mathcal{C}$
Figure 1: Overview of TextOCVP. (a) Our model first parses a reference frame $\textbf{X}_{1}$ into its object components $\textbf{S}_{1}$. (b) Our TextOCVP predictor jointly models object dynamics and interactions while integrating textual guidance, generating future object states and video frames that align with the provided textual instructions.
1 Introduction
Understanding and reasoning about the real world is essential for enabling autonomous systems to better comprehend their surroundings, predict future events, and adapt their actions accordingly. Humans achieve these capabilities by perceiving the environment as a structured composition of individual objects that interact and evolve dynamically over time (Kahneman et al., 1992). Neural networks equipped with such compositional inductive biases have shown the ability to learn structured object-centric representations of the world, which enable desirable properties, such as out-of-distribution generalization (Dittadi et al., 2022), compositionality (Greff et al., 2020), or sample efficiency (Mosbach et al., 2024).
Recent advances in unsupervised object-centric representation learning have progressed from extracting object representations in synthetic images (Locatello et al., 2020; Lin et al., 2020b) to modeling objects in video (Kipf et al., 2022; Singh et al., 2022) and scaling to real-world scenes (Seitzer et al., 2023; Zadaianchuk et al., 2024). These developments have enabled object-level dynamics modeling for future prediction and planning. Notably, models like SlotFormer (Wu et al., 2023) or OCVP (Villar-Corrales et al., 2023) introduced object-centric world models that explicitly model spatio-temporal relationships between objects, shifting away from image-level approaches that ignore scene compositionality. Despite these advancements, current approaches struggle with complex object appearances and dynamics, and lack the ability to incorporate external guidance, thus limiting their effectiveness as world models in robotic applications.
To address these challenges, we propose TextOCVP, a novel object-centric model for image-to-video generation with language guidance, illustrated in Fig. 1. Given a reference image and text instruction, TextOCVP extracts object representations and predicts their evolution using a novel text-conditioned object-centric transformer. This module forecasts future object representations by modeling object dynamics and interactions, while also incorporating textual information through a text-to-slot attention mechanism. By jointly modeling spatio-temporal object relationships and integrating textual guidance, TextOCVP generates future object states and video frames that align with the provided instructions.
We evaluate our approach through extensive experiments focusing on image-to-video generation tasks with varying levels of complexity. Our results show that TextOCVP outperforms existing text-conditioned methods by effectively leveraging object-centric representations. This demonstrates the significant advantage of incorporating textual guidance into structured object-centric representations, particularly for scenes featuring multiple moving objects.
Through an in-depth model analysis, we demonstrate the importance of our proposed object-centric latent space, which allows TextOCVP to generate video sequences that closely align with language instructions. We further demonstrate how TextOCVP adapts its generated video sequences based on different captions by routing the text information to the relevant object representations.
@@ 1 Introduction
  • We propose TextOCVP, a novel text-driven image-to-video generation model, featuring a text-conditioned object-centric predictor that integrates textual guidance via a text-to-slot attention mechanism.
  • Through extensive evaluation, we show that TextOCVP outperforms existing text-conditioned models by leveraging object-centric representations.
  • We demonstrate that TextOCVP is controllable, seamlessly adapting to diverse textual instructions.
@@ 2.1 Object-Centric Learning
Representation learning, the ability to extract meaningful features from data, often improves model performance by enhancing its understanding of the input space (Bengio et al., 2013). Object-centric representation learning focuses on simultaneously learning representations that characterize individual objects within an input image or video.
Recently, object-centric models have progressed from learning object representations from synthetic images (Greff et al., 2019; Lin et al., 2020a; Villar-Corrales & Behnke, 2022; Locatello et al., 2020) to videos (Kipf et al., 2022; Singh et al., 2022; Elsayed et al., 2022), and real-world scenes (Seitzer et al., 2023; Aydemir et al., 2023; Zadaianchuk et al., 2024; Kakogeorgiou et al., 2024). The learned object representations benefit downstream tasks, such as reinforcement learning (Mosbach et al., 2024) or visual-question answering (Wu et al., 2023), among others.
2.2 Video Prediction and Generation
@@
Future frame video prediction (VP) is the task of forecasting the upcoming $T$ video frames conditioned on the preceding $C$ seed frames (Oprea et al., 2020). When the number of seed frames is $C=1$, this task is often referred to as image-to-video generation. Several methods have been proposed to address this challenge, leveraging 2D convolutions (Gao et al., 2022; Chiu et al., 2020), 3D convolutions (Vondrick et al., 2016; Tulyakov et al., 2018), recurrent neural networks (RNNs) (Denton & Fergus, 2018; Villar-Corrales et al., 2022; Wang et al., 2022; Guen & Thome, 2020; Franceschi et al., 2020), Transformers (Rakhimov et al., 2021; Ye & Bilodeau, 2022, 2023), or diffusion models (Höppe et al., 2022; Ho et al., 2022).
2.2.1 Object-Centric Video Prediction
@@
Object-centric VP presents a structured approach that explicitly models the dynamics and interactions of individual objects to forecast future video frames. These methods typically involve three main steps: decomposing seed frames into object representations, forecasting future object states using a dynamics model, and rendering video frames from the predicted object representations. Various approaches have been proposed for this task, using different architectural designs such as RNNs (Creswell et al., 2021; Zoran et al., 2021; Nguyen et al., 2024) or transformers (Wu et al., 2021, 2023; Villar-Corrales et al., 2023; Song et al., 2023; Daniel & Tamar, 2024). Despite promising results, these models are currently limited to simple deterministic datasets or rely on action-conditioning (Mosbach et al., 2024) or inferred latent vectors (Villar-Corrales & Behnke, 2025). In contrast, our proposed model forecasts future video frames conditioned on past object slots and textual descriptions.
2.2.2 Text-Conditioned Video Prediction
This category of VP models leverages text descriptions to provide appearance, motion and action cues that guide the generation of future frames. This task was first proposed by Hu et al. (2022), who utilized a VQ-VAE to encode the images into visual token representations, and modeled the image dynamics with an axial transformer that jointly processes such tokens along with text descriptions. TVP (Song et al., 2024) leverages RNNs to learn text representations, which condition a GAN-based framework in order to generate videos from a single frame. MMVG (Fu et al., 2023) combines a VQ-GAN (Esser et al., 2021) with a masked transformer predictor. More recently, several methods leverage diffusion models with text conditioning (Gu et al., 2024; Ni et al., 2023; Chen et al., 2023).
@@ 2.2.2 Text-Conditioned Video Prediction
Concurrently with our work, Wang et al. (2024) combine object-centric learning with text-conditional image-to-video generation. The authors propose an autoregressive diffusion model conditioned on object slots and text descriptions and evaluate on simple datasets. In contrast, we explicitly model the object dynamics with an autoregressive transformer, and evaluate our approach on more complex robotics simulations.
@@ 3 Method
We propose TextOCVP – a novel object-centric model for text-conditioned image-to-video generation. Given an initial reference image $\textbf{X}_{1}$ and a text caption $\mathcal{C}$, TextOCVP generates the subsequent $T$ video frames $\hat{\mathbf{X}}_{2:{T+1}}$, which maintain a similar appearance and structural composition as the reference image, and follow the motion described in the text caption.
TextOCVP, which is illustrated in Fig. 2, implements an object-centric approach, in which the reference frame $\textbf{X}_{1}$ is first decomposed with a scene parsing module (Sec. 3.1) into a set of $N_{\mathbf{S}}$ $D$-dimensional object representations called slots $\textbf{S}_{1}\in{\mathbb{R}}^{N_{\mathbf{S}}\times D}$, where each slot represents a single object in the image. The object slots are fed to a text-conditioned transformer predictor (Sec. 3.2), which jointly models their spatio-temporal relations, and incorporates the textual information from the caption $\mathcal{C}$ as a guidance for predicting the future object slots $\hat{\mathbf{S}}_{2:T+1}$. Finally, the predicted object slots are decoded in order to render object representations and video frames (Sec. 3.3).
We propose two different TextOCVP variants, which differ in the underlying object-centric decomposition modules. Specifically, $\text{TextOCVP}_{\text{SAVi}}$ leverages SAVi (Kipf et al., 2022), whereas $\text{TextOCVP}_{\text{DINO}}$ extends the DINOSAUR (Seitzer et al., 2023) framework for recursive object-centric video decomposition and video rendering.
Figure 2: Overview of TextOCVP. TextOCVP parses the reference frame $\textbf{X}_{1}$ into object representations $\textbf{S}_{1}$. The text-conditioned object-centric predictor models object dynamics and interactions, incorporating information from the description $\mathbf{C}$ to predict future object states and frames.
@@ 3.1 Scene Parsing
The scene parsing module decomposes a video sequence $\mathbf{X}_{1:{\tau}}$ into a set of permutation-invariant object representations called slots $\mathbf{S}_{1:\tau}=(\textbf{S}_{1},...,\textbf{S}_{\tau})$, with $\textbf{S}_{t}=(\textbf{s}_{t}^{1},...,\textbf{s}_{t}^{N_{\mathbf{S}}})$, where each slot s represents a single object.
For learning the object representations, we leverage the object-centric video decomposition framework proposed by Kipf et al. (2022). At time step $t$, the corresponding input frame $\textbf{X}_{t}$ is encoded with a feature extractor module into a set of $D_{h}$-dimensional feature maps $\textbf{h}_{t}\in{\mathbb{R}}^{L\times D_{h}}$ representing $L$ spatial locations. The feature extractor is convolutional neural network in our $\text{TextOCVP}_{\text{SAVi}}$ variant and a DINO-pretrained vision transformer (Caron et al., 2021) in $\text{TextOCVP}_{\text{DINO}}$. These feature maps are processed with a Slot Attention (Locatello et al., 2020) corrector, which updates the previous slots $\textbf{S}_{t-1}$ based on visual features from the current frame following an iterative attention mechanism. Namely, Slot Attention performs cross-attention, with the attention weights normalized over the slot dimension, thus encouraging the slots to compete to represent parts of the input. It then updates the slots using a Gated Recurrent Unit (Cho et al., 2014) (GRU). Formally, Slot Attention updates the previous slots $\textbf{S}_{t-1}$ by:
@@
$\textbf{S}_{t}=\text{GRU}(W_{t}v(\textbf{h}),\textbf{S}_{t-1})\;\text{with}\;W_{i,j}=\frac{\textbf{A}_{i,j}}{\sum_{l=1}^{L}\textbf{A}_{i,l}},$ (2)
where $k,q$ and $v$ are learned linear projections that map input features and slots into a common dimension. The final output of this module is the set of slots $\textbf{S}_{t}$, representing the objects of the input frame.
@@ 3.2 Text-Conditioned Object-Centric Predictor
To condition the prediction process, the text description $\mathcal{C}$ is encoded into text token embeddings $\mathbf{C}$ employing an encoder-only transformer module. We experiment with different variants of this module, including a vanilla transformer encoder (Vaswani et al., 2017) or a pretrained T5 (Raffel et al., 2020) module.
At time step $t$, the predictor receives as input the corresponding text embeddings $\mathbf{C}$, as well as the previous object slots $\mathbf{S}_{1:t}$, which are initially mapped via an MLP to the predictor token dimensionality $D_{\text{Pred}}$. Additionally, these tokens are augmented with a temporal positional encoding, which applies the same sinusoidal positional embedding to all tokens from the same time-step, thus preserving the inherent permutation-equivariance of the objects.
Each layer of our predictor module mirrors the transformer decoder architecture (Vaswani et al., 2017). First, a self-attention layer enables every slot to attend to all other object representations in the sequence, modeling the spatio-temporal relations between objects. Subsequently, a text-to-slot cross-attention layer enhances the slot representations by incorporating important features from the text embeddings, such as motion or appearance information. Finally, an MLP is independently applied for each token. This process is repeated in every predictor layer, resulting in the predicted object slots of the subsequent time step $\hat{\textbf{S}}_{t+1}$. Furthermore, we apply a residual connection from $\textbf{S}_{t}$ to $\hat{\textbf{S}}_{t+1}$, which improves the temporal consistency of the predictions. This process is repeated autoregressively to obtain slot predictions for $T$ subsequent time steps.

### v2

# TextOCVP: Object-Centric Video Prediction with Language Guidance

@@ Abstract
Understanding and forecasting future scene states is critical for autonomous agents to plan and act effectively in complex environments. Object-centric models, with structured latent spaces, have shown promise in modeling object dynamics and predicting future scene states, but often struggle to scale beyond simple synthetic datasets and to integrate external guidance, limiting their applicability in robotics. To address these limitations, we propose TextOCVP, an object-centric model for video prediction guided by textual descriptions. TextOCVP parses an observed scene into object representations, called slots, and utilizes a text-conditioned transformer predictor to forecast future object states and video frames. Our approach jointly models object dynamics and interactions while incorporating textual guidance, enabling accurate and controllable predictions. TextOCVP’s structured latent space offers a more precise control of the forecasting process, outperforming several video prediction baselines on two datasets. Additionally, we show that structured object-centric representations provide superior robustness to novel scene configurations, as well as improved controllability and interpretability, enabling more precise and understandable predictions. Videos and code are available at https://play-slot.github.io/TextOCVP.
1 Introduction
Understanding and reasoning about the environment is essential for enabling autonomous systems to better comprehend their surroundings, predict future events, and adapt their actions accordingly. Humans achieve these capabilities by perceiving the environment as a structured composition of individual objects that interact and evolve dynamically over time (Kahneman, Treisman, and Gibbs 1992). Neural networks equipped with such compositional inductive biases have shown the ability to learn structured object-centric representations of the world, which enable desirable properties, such as out-of-distribution generalization (Dittadi et al. 2022), compositionality (Greff, Van Steenkiste, and Schmidhuber 2020), or sample efficiency (Mosbach et al. 2025).
Recent advances in unsupervised object-centric representation learning have progressed from extracting object representations in synthetic images (Locatello et al. 2020; Lin et al. 2020) to modeling objects in video (Kipf et al. 2022; Singh, Wu, and Ahn 2022) and scaling to real-world scenes (Seitzer et al. 2023; Zadaianchuk, Seitzer, and Martius 2024). These developments have enabled object-level dynamics modeling for future prediction and planning. Notably, approaches like SlotFormer (Wu et al. 2023) or OCVP (Villar-Corrales, Wahdan, and Behnke 2023) introduced object-centric prediction models that explicitly model spatio-temporal relationships between objects, shifting away from image-level approaches that ignore scene compositionality. Despite these advancements, current object-centric methods struggle with complex object appearances and dynamics, and lack mechanisms to incorporate external guidance, thus limiting their scalability and broader applicability.
Refer to captiona) Object-Centric ParsingLearned ObjectsRefer to captionb) Text-Conditioned Video Prediction
@@ 1 Introduction
‘Put the green block
in the red bowl.’
Figure 1: Overview of TextOCVP. (a) Our model parses a reference frame $\textbf{X}_{1}$ into its object components $\textbf{S}_{1}$. (b) Our TextOCVP predictor jointly models object dynamics and interactions guided by text, generating future object states and frames that align with the provided textual instructions.
To address these challenges, we propose TextOCVP, a novel object-centric model for video prediction guided by textual instructions, illustrated in Fig. 1. Given a reference image and text instruction, TextOCVP extracts object representations and predicts their evolution using a text-conditioned object-centric transformer. This predictor forecasts future object states by explicitly modeling their dynamics and interactions over time, while integrating textual information via a text-to-slot attention mechanism. By jointly modeling spatio-temporal object relationships and incorporating textual guidance, TextOCVP predicts future object states and frames aligned with the input instruction.
We evaluate our approach for text-conditioned video prediction through extensive experiments on two distinct datasets. Our results show that TextOCVP outperforms existing text-conditioned video prediction methods by effectively leveraging structured object-centric representations, particularly on scenes featuring multiple moving objects.
To verify the effectiveness of object-centric representations for text-guided video prediction, we conduct an in-depth analysis of TextOCVP. Specifically, we demonstrate that its structured latent space enables accurate and controllable video prediction by aligning language instructions with the corresponding objects, outperforming baselines that rely on holistic scene representations. Beyond generation quality, TextOCVP offers improved interpretability and exhibits strong robustness to novel scene configurations, including varying numbers of objects or previously unseen colors.
@@
  • We propose TextOCVP, a text-guided video prediction model, featuring a text-conditioned object-centric predictor that integrates textual guidance into the prediction process via a text-to-slot attention mechanism.
  • Through extensive evaluations, we show that TextOCVP outperforms existing text-conditioned video prediction models by leveraging object-centric representations.
  • We demonstrate that TextOCVP is controllable, seamlessly adapting to diverse textual instructions, while exhibiting interpretability and robustness to novel setups.
@@ 2.1 Object-Centric Learning
Representation learning, the ability to extract meaningful features from data, often improves model performance by enhancing its understanding of the input space (Bengio, Courville, and Vincent 2013). Object-centric representation methods aim to parse an image or video into a set of object components in an unsupervised manner. These objects are typically represented as unconstrained embeddings (called slots) (Locatello et al. 2020; Kipf et al. 2022; Singh, Wu, and Ahn 2022), patch-based representations (Lin et al. 2020), factored latent vectors (Greff et al. 2019), or explicit object prototypes (Villar-Corrales and Behnke 2022; Monnier, Groueix, and Aubry 2020). These methods have demonstrated promise in learning object representations across diverse domains, ranging from synthetic images (Locatello et al. 2020; Lin et al. 2020) to videos (Kipf et al. 2022; Singh, Wu, and Ahn 2022; Elsayed et al. 2022), and real-world scenes (Seitzer et al. 2023; Zadaianchuk, Seitzer, and Martius 2024). The learned object representations benefit downstream tasks, such as reinforcement learning (Mosbach et al. 2025) or visual-question answering (Wu et al. 2023).
2.2 Video Prediction
@@
Video prediction (VP) is the task of forecasting the upcoming $T$ video frames conditioned on the preceding $C$ seed frames (Oprea et al. 2020). Several methods have been proposed to address this challenge, leveraging 2D convolutions (Gao et al. 2022; Chiu, Adeli, and Niebles 2020), 3D convolutions (Tulyakov et al. 2018), recurrent neural networks (RNNs) (Denton and Fergus 2018; Villar-Corrales et al. 2022; Wang et al. 2022), Transformers (Rakhimov et al. 2021; Ye and Bilodeau 2022), or diffusion models (Höppe et al. 2022; Ho et al. 2022).
Object-Centric Video Prediction
@@
Object-centric VP presents a structured approach that explicitly models the dynamics and interactions of individual objects to forecast future video frames. These methods typically involve three main steps: parsing seed frames into object representations, forecasting future object states using a dynamics model, and rendering video frames from the predicted object representations. Various approaches have addressed this task using different architectural priors, such as RNNs (Creswell et al. 2021; Nguyen et al. 2024) or transformers (Wu, Yoon, and Ahn 2021; Wu et al. 2023; Villar-Corrales, Wahdan, and Behnke 2023). Despite promising results, these models are limited to simple deterministic datasets or rely on action-conditioning (Mosbach et al. 2025; Villar-Corrales and Behnke 2025). In contrast, our model forecasts future frames conditioned on past object slots and text descriptions.
Text-Conditioned Video Prediction
Text-conditioned VP models leverage text descriptions to provide appearance, motion and action cues that guide the generation of future frames. This task was first proposed by Hu, Luo, and Chen (2022), who utilized a VQ-VAE to encode images into visual token representations, and modeled the scene dynamics with an axial transformer to jointly process visual tokens with text descriptions. Similarly, approaches like TVP (Song et al. 2024) and MMVG (Fu et al. 2023) address this task using RNNs or masked transformers, respectively. More recently, several methods leverage diffusion models for text-guided VP (Blattmann et al. 2023; Ni et al. 2023; Chen et al. 2024; Xing et al. 2024; Chen et al. 2023; Gu et al. 2024). These approaches encode video frames into discrete token sequences via pretrained quantized autoencoders, and leverage pretrained language models to guide the diffusion-based generation with text features. To further improve temporal coherence and semantic alignment, several diffusion-based methods leverage specialized conditioning strategies (Xing et al. 2024; Chen et al. 2023) or attention mechanisms (Gu et al. 2024), among others. While effective, these models operate on holistic or spatial representations, and require large-scale compute and data. In contrast, TextOCVP adopts a more structured and efficient approach, explicitly modeling object dynamics using slot representations, where each object in the scene is represented by a distinct embedding.
@@
Concurrently with our work, Wang et al. (2025) and Jeong et al. (2025) combine object-centric learning with autoregressive diffusion models and transformers for text-guided video prediction on simple synthetic datasets. In contrast, we model object dynamics with an autoregressive transformer, evaluate on more complex robotics simulations, and perform an in-depth analysis of the properties of object-centric representations for text-conditioned video prediction.
@@ 3 Method
We propose TextOCVP, a novel object-centric model for text-conditioned video prediction. Given an initial reference image $\textbf{X}_{1}$ and a text caption $\mathcal{C}$, TextOCVP generates the $T$ subsequent video frames $\hat{\mathbf{X}}_{2:{T+1}}$, which maintain a similar appearance and structural composition as the reference image, and follow the motion described in the text caption.
TextOCVP, which is illustrated in Fig. 2, implements an object-centric approach, in which the reference frame $\textbf{X}_{1}$ is first decomposed with a scene parsing module (Sec. 3.1) into a set of $N_{\mathbf{S}}$ $D$-dimensional object representations called slots $\textbf{S}_{1}\in{\mathbb{R}}^{N_{\mathbf{S}}\times D}$, where each slot represents a single object in the image. The object slots are fed to a text-conditioned transformer predictor (Sec. 3.2), which jointly models their spatio-temporal relations, and incorporates the textual information from the caption $\mathcal{C}$ as guidance for predicting the future object slots $\hat{\mathbf{S}}_{2:T+1}$. Finally, the predicted slots are decoded to render future video frames (Sec. 3.3).
We propose two different TextOCVP variants, which differ in the underlying object-centric decomposition modules. Specifically, $\text{TextOCVP}_{\text{SAVi}}$ leverages SAVi (Kipf et al. 2022), whereas $\text{TextOCVP}_{\text{DINO}}$ extends the recent DINOSAUR (Seitzer et al. 2023) framework for recursive object-centric video decomposition and video rendering.
@@ 3.1 Scene Parsing
The scene parsing module decomposes a video sequence $\mathbf{X}_{1:{\tau}}$ into a set of permutation-invariant object representations called slots $\mathbf{S}_{1:\tau}=(\textbf{S}_{1},...,\textbf{S}_{\tau})$, with $\textbf{S}_{t}=(\textbf{s}_{t}^{1},...,\textbf{s}_{t}^{N_{\mathbf{S}}})$, where each slot $\textbf{s}\in{\mathbb{R}}^{D}$ represents a single object. For scene parsing, we adopt the recursive object-centric video decomposition framework from Kipf et al. (2022).
At time step $t$, the corresponding input frame $\textbf{X}_{t}$ is encoded with a feature extractor module into a set of $D_{h}$-dimensional feature maps $\textbf{h}_{t}\in{\mathbb{R}}^{L\times D_{h}}$ representing $L$ spatial locations. The feature extractor is a convolutional neural network in our $\text{TextOCVP}_{\text{SAVi}}$ variant and a DINO-pretrained vision transformer (Caron et al. 2021) in $\text{TextOCVP}_{\text{DINO}}$. These feature maps are processed with Slot Attention (Locatello et al. 2020), which updates the previous slots $\textbf{S}_{t-1}$ based on visual features from the current frame following an iterative attention mechanism. Namely, Slot Attention performs cross-attention, with the attention weights normalized over the slot dimension, thus encouraging the slots to compete to represent parts of the input. It then updates the slots using a Gated Recurrent Unit (Cho et al. 2014) (GRU). Formally, Slot Attention updates the previous slots $\textbf{S}_{t-1}$ by:
@@
$\textbf{S}_{t}=\text{GRU}(W_{t}v(\textbf{h}_{t}),\textbf{S}_{t-1})\;\text{with}\;W_{i,j}=\frac{\textbf{A}_{i,j}}{\sum_{l=1}^{L}\textbf{A}_{i,l}},$ (2)
where $k,q$ and $v$ are learned linear projections that map input features and slots into a common dimension. The output of this module is a set of slots $\textbf{S}_{t}$ that represents the objects of the input frame.
Figure 2: Overview of TextOCVP. TextOCVP parses the reference frame $\textbf{X}_{1}$ into object representations $\textbf{S}_{1}$. The text-conditioned object-centric predictor models object dynamics and interactions, incorporating information from the description $\mathbf{C}$ to predict future object states $\hat{\mathbf{S}}_{2:T+1}$, which can be used to render future video frames $\hat{\mathbf{X}}_{2:{T+1}}$.
@@ 3.2 Text-Conditioned Object-Centric Predictor
To condition the prediction process, the text description $\mathcal{C}$ is encoded into text token embeddings $\mathbf{C}$ employing an encoder-only transformer module. We experiment with different variants of this module, including a vanilla transformer encoder (Vaswani et al. 2017) or a pretrained T5 (Raffel et al. 2020) module.
At time step $t$, the predictor receives as input the corresponding text embeddings $\mathbf{C}$, as well as the previous object slots $\mathbf{S}_{1:t}$, which are initially mapped via an MLP to the predictor token dimensionality. Additionally, these tokens are augmented with a temporal positional encoding, which applies the same sinusoidal positional embedding to all tokens from the same time step, thus preserving the inherent permutation-equivariance of the objects.
Each layer of our predictor module mirrors the transformer decoder architecture (Vaswani et al. 2017). First, a self-attention layer enables every slot to attend to all other object representations in the sequence, modeling the spatio-temporal relations between objects. Subsequently, a text-to-slot cross-attention layer enhances the slot representations by incorporating relevant features from the text embeddings, guiding the prediction process to align with the motion and dynamics described in the textual caption. Finally, an MLP is applied for each token. This process is repeated in every predictor layer, resulting in the predicted object slots of the subsequent time step $\hat{\textbf{S}}_{t+1}$. Furthermore, we apply a residual connection from $\textbf{S}_{t}$ to $\hat{\textbf{S}}_{t+1}$, which improves the prediction temporal consistency. This process is repeated autoregressively to obtain slot predictions for $T$ subsequent time steps.